# Quantum Singular Value Transformation for Solving the Time-Dependent Maxwell’s Equations

### v1

@@ 1 Introduction
Linear systems of equations play an important role in many areas of science and technology. For example, in biology they arise in modeling metabolic networks, population dynamics, and biochemical reaction pathways [1, 2, 3]. In machine learning, training algorithms such as linear regression or support vector machines often require solving systems of equations derived from optimization objectives [4]. In computational physics and engineering, discretizing partial differential equations using finite difference or finite element methods reduces the problem to solving large, sparse linear systems [5]. In fluid dynamics, the discretization and linearization of the Navier–Stokes equations during numerical simulation also necessitates robust and efficient solvers [6]. The ubiquity of linear systems across scientific disciplines highlights the fundamental importance of developing scalable algorithms for their solution.
@@
In the past decade, significant attention has been given to the possibility of solving linear systems on quantum computers. Classically solving an N $\times$ N linear system (N equations for N unknowns) scales polynomially in N. In contrast, Harrow-Hassidim-Lloyd (HHL) [7] introduced a quantum algorithm that scales logarithmically in N, suggesting that quantum computers may provide an exponential speedup for certain linear system problems. More recently, the Variational Quantum Linear Solver (VQLS) [8] has been proposed as a hybrid quantum-classical alternative that is more suitable for near-term quantum devices. VQLS formulates the solution of $Ax=b$ as a variational optimization problem, where a parameterized quantum circuit prepares an approximate solution state $|x\rangle$ by minimizing a cost function that encodes the residual error in the linear system. This method avoids deep quantum circuits and phase estimation, making it more practical under realistic noise and coherence constraints. An alternative, fully quantum paradigm is the Quantum Singular Value Transformation (QSVT) [9], which provides a powerful and modular framework for implementing matrix functions on quantum states. QSVT leverages block-encoded matrices and polynomial approximations to perform transformations directly on the singular values of an operator. By constructing a polynomial approximation to the reciprocal function $f(x)\approx 1/x$ over the spectrum of $A$, QSVT enables the implementation of an inverse transformation $A^{-1}$ as a quantum circuit. This approach generalizes and improves upon earlier linear system solvers by avoiding phase estimation and offering provable bounds on approximation error and complexity, while remaining compatible with fault-tolerant quantum computation.
@@ 3 Methods
One of the most common ways to represent a linear system is in matrix-vector form: given an invertible matrix $A\in\mathbb{R}^{N\times N}$ and a vector $\vec{b}$, the aim is to solve $A\vec{x}=\vec{b}$, which requires computing $\vec{x}=A^{-1}\vec{b}$. Using QSVT, the inverse operation is approximated by applying a polynomial transformation to the singular values of $A$. In practice, this means constructing a polynomial $P(x)\approx\frac{1}{x}$, implemented through a sequence of controlled unitaries parameterized by phase angles. This section deals with the construction of such polynomial approximations using Quantum Singular Value Transformation (QSVT), detailing the design of the phase factors and the implementation of the corresponding quantum circuits for approximating the inverse of $A$.
@@ 3.1 Padé scheme
The system can be written in matrix form
@@ 3.4 Polynomial design
To solve Maxwell’s equations a 21st-order polynomial is used, resulting in a sum of even and odd polynomials of degrees 10 and 11, respectively. The reason for using an even and odd polynomials is explained in Sec. 3.5.

### v2

@@ 1 Introduction
The fundamental principles of classical electromagnetism are encapsulated by Maxwell’s equations combined with the Lorentz force law, which serve as the theoretical basis for many technologies including optics, wireless communication systems, and electronic circuitry [1, 2, 3]. Due to their inherent nonlinearity and hyperbolic nature, obtaining exact analytical solutions to these equations is challenging, particularly in multidimensional scenarios. As a result, computational approaches such as the finite-difference time-domain (FDTD) technique [4], finite-element methods [5], and finite-volume schemes [6] have become essential tools for producing approximate numerical solutions.
@@
These numerical methods often rely on solving large systems of linear equations, which arise naturally from the discretization of Maxwell’s equations and many other partial differential equations. Linear systems of equations play an important role in many areas of science and technology. For example, in biology they arise in modeling metabolic networks, population dynamics, and biochemical reaction pathways [7, 8, 9]. In machine learning, training algorithms such as linear regression or support vector machines often require solving systems of equations derived from optimization objectives [10]. In computational physics and engineering, discretizing partial differential equations using finite difference or finite element methods reduces the problem to solving large, sparse linear systems [11]. In fluid dynamics, the discretization and linearization of the Navier–Stokes equations during numerical simulation also necessitates robust and efficient solvers [12]. The ubiquity of linear systems across scientific disciplines highlights the fundamental importance of developing scalable algorithms for their solution.
@@
In the past decade, significant attention has been given to the possibility of solving linear systems on quantum computers. Classically solving an N $\times$ N linear system (N equations for N unknowns) scales polynomially in N. In contrast, Harrow-Hassidim-Lloyd (HHL) [13] introduced a quantum algorithm that scales logarithmically in N, suggesting that quantum computers may provide an exponential speedup for certain linear system problems. More recently, the Variational Quantum Linear Solver (VQLS) [14] has been proposed as a hybrid quantum-classical alternative that is more suitable for near-term quantum devices. VQLS formulates the solution of $Ax=b$ as a variational optimization problem, where a parameterized quantum circuit prepares an approximate solution state $|x\rangle$ by minimizing a cost function that encodes the residual error in the linear system. This method avoids deep quantum circuits and phase estimation, making it more practical under realistic noise and coherence constraints. However, VQLS relies on repeated ansatz optimization, which can be computationally expensive and less practical for solving time-dependent PDEs that require solutions at multiple time steps. An alternative, fully quantum paradigm is the Quantum Singular Value Transformation (QSVT) [15], which provides a powerful and modular framework for implementing matrix functions on quantum states. QSVT leverages block-encoded matrices and polynomial approximations to perform transformations directly on the singular values of an operator. By constructing a polynomial approximation to the reciprocal function $f(x)\approx 1/x$ over the spectrum of $A$, QSVT enables the implementation of an inverse transformation $A^{-1}$ as a quantum circuit. This approach generalizes and improves upon earlier linear system solvers by avoiding phase estimation and offering provable bounds on approximation error and complexity, while remaining compatible with fault-tolerant quantum computation.
@@ 3 Methods
One of the most common ways to represent a linear system is in matrix-vector form: given an invertible matrix $A\in\mathbb{R}^{N\times N}$ and a vector $\vec{b}$, the aim is to solve $A\vec{x}=\vec{b}$, which requires computing $\vec{x}=A^{-1}\vec{b}$. Using QSVT, the inverse operation is approximated by applying a polynomial transformation to the singular values of $A$. In practice, this means constructing a polynomial $P(t)\approx\frac{1}{t}$, implemented through a sequence of controlled unitaries parameterized by phase angles. This section deals with the construction of such polynomial approximations using Quantum Singular Value Transformation (QSVT), detailing the design of the phase factors and the implementation of the corresponding quantum circuits for approximating the inverse of $A$.
@@ 3.1 Padé scheme
Writing this discrete equation at every grid point results in the need to solve a system of linear algebraic equations to compute the discrete derivative, and it can be represented in the following form
@@ 3.4 Polynomial design
To solve Maxwell’s equations a 21st-order polynomial is used, resulting in a sum of even and odd polynomials of degrees 10 and 11, respectively. This order was chosen because it required relatively low computational effort while delivering high-fidelity results. The reason for using an even and odd polynomials is explained in Sec. 3.5.